A small manufacturing company has a demand of 800 units of its products per month. The process is completed in four stages, which produce the following average percentages of good-quality units: Stage 1 Stage 2 Stage 3 Stage 4 88% 92% 81% 95% How many units of components does it need to buy to satisfy demand?

a. 701 – 950

b. 1201 – 1450

c. More than 1450

d. Less than or equal to 700

e. 951 – 1200
Answer #1

Most appropriate response: (b) 1201 - 1450

Justification:

Let the components purchased be 'x'
Therefore,
Stage Input Efficiency Output
Stage 1 x 88% 0.88x
Stage 2 0.88x 92% (0.88 * 0.92)x = 0.8096x
Stage 3 0.8096x 81% (0.8096 * 0.81)x = 0.6558x
Stage 4 0.6558x 95% (0.6558 * 0.95)x = 0.623x
Therefore,
0.623x =800 units
x = 1284.1091 i.e option (b)